Help me with this in a hurry please? I need it explained all problems algebraically please?



INVESTMENT CLUBS: Part of $8000.00 was invested by an investment club at 10% interest and the rest at 12%. If the annual income from these investments is $900.00, how much was invested at each rate?
Thank you so much!!

Answer by venugopalramana(3286)

Part of $8000.00 was invested by an investment club at 10% interest
LET THIS INVESTMENT BE X...SO INTERST ON THIS IS..I=PNR/100=X*1*10/100=10X/100
and the rest at 12%.
REST OF THE MONEY IS (8000-X)...SO INTEREST ON THIS IS =(8000-X)*1*12/100
If the annual income from these investments is $900.00, how much was invested at each rate?
TOTAL INCOME =10X/100 + (8000-X)*12/10=900
(10X+96000-12X)/100=900
-2X+96000=90000
2X=96000-90000=6000
X=3000
HENCE 3000 IS INVESTED AT 10% AND REST 5000 INVESTED AT 12%